A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
The answer I got is 5 hours

Answer by stanbon(48558)

A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
The answer I got is 5 hours
----------------
It took the plane one hour to fly 560-480= 80 miles
distance/rate = 480/80 = 6 hours for the 2nd part of the journey.
===========
Cheers,
Stan H.